A recent downturn in economy is causing a man's 401(k) to lose value at a linear rate. On january 1 the account was worth $120,000 and on August 1 it was worth $96,000. If the current trend continues,on December 1 it will be worth how much?
Answer by checkley77(12844)

Using the line equation for this situation we get the 2 points:
(0, 120,000) (8, 96,000)
Y=mX+b where m=slope & b=Y intercept.
First we find the slope:
Slope=(Y2-Y1)/(X2-X1)
Slope=(96,000-120,000)/(8-0)
Slope=(-24,000)/8=-3,000
12000=-3,000*0+b
b=120000 the Y intercept.
y=-3,000x+120,000 is the line equation.
Y=-3,000*11+120,000
Y=-33,000+120,000
y=87,000 is the value of the account on Dec. 1.